Graph parabola show the vertex and at least two points on either side of the vertex.
y = x^2 -4x +3

Answer by richwmiller(17219)

h=-b/2a
h=-(-4)/2
h=2
y = 2^2 -4*2 +3
y=4-8+3
y=-1
the vertex=(2,-1)
x=1 and x=3
are on either side of the vertex
y=1^2-4+3
y=0
(1,0)
y=3^2-4(3)+3
y=9-12+3
y=0
(3,0)